Actions
Ops/Comm Scheduling meeting : 11/7/2017¶
Location : CR21 at 14:00 pm
Phone access : New phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517
- Actions
- Hall Manager - water system change-over procedure (stands) - have prescription and draft document (Colin/Josef), stands
- Ajit and Henry have worked through the issues list: C&M and Beamline issues have been identified and assigned.
Steve Boyd will go through the Other issues with Colin Whyte next week - Colin and I will go through this on Wednesday
- DC update
- Hydrogen system update and plans
- Cycle 2017 Issues: To be gone through tomorrow
- Status of Support for Technical Endeavours And Knowledge (CWhyte)
- AOB
- Last meeting : Tuesday June 4 2017
- Next meeting : Tuesday July 18 2017
Minutes¶
- Actions
- Hall Manager - water system change-over procedure (stands) - have prescription and draft document (Colin/Josef), stands
JBoehm has made a dry-run, working through the document. Procedure seems OK, but would like larger flags on the valves. Minor changes then publish document. - Cycle 2017 Issues: Ajit and Henry have worked through the issues list: C&M and Beamline issues have been identified and assigned.
Steve Boyd will go through the Other issues with Colin Whyte next week - Colin and I will go through this on Wednesday
Also need to check the paper logbook for issues listed there.
- Hall Manager - water system change-over procedure (stands) - have prescription and draft document (Colin/Josef), stands
- DC update
Steve and Trevor down last week. Checked through the interlocks on the Cooling Channel magnets, then powered all into shorts, then 10 A into the SS's. All OK apart from one interlock on the cooling water. PSUs left isolated. Still need to correct the solenoid/flip configuration. The heaters were left off afterwards.
Ed and Joe got (finally) the target control PC running in the MLCR. Moved it across to R78: discovered the VME crate and the Target Controller were both broken. Have brought the spare controller down from Sheffield, are finally running in R78 and taking data. The carrier board in the IOC was changed - we still have spares left.
Tracker cryostats warm and being pumped out. Purge cycles being done less often to reduces stress on cassettes. The Tracker vacuum readout has been fixed by correcting the brainbox configuration. Spare scroll pump on order.
Air-conditioning precarious: RR2 is stable (on one of two) but the Hall has two units down. Craig has chased Paul Masterson to get the engineers in. JBoehm reported that he had found high pressure (1150 vs 1060 mbar) in the FC when he had come in on Monday, A/C 5 was off and the Hall was very warm (29°C). Cryo-cooler was fine so believes this is due to heat leaking into the FC - need to be aware of aircon failures during running. The A/C units (and Hall temperatures) are in the Alarm Handler. The temperature readouts from the environment monitoring appear to be stalled. Note that sub-system Alarm Handlers (e.g. LH2) need to include their ancillary services. JBoehm was concerned about the number of alarms going off: how do we ensure shifters actually respond to an LH2 alarm? Possibly it could use a different sound?
- Hydrogen system update and plans
JBoehm working on the documentation.
The production turret has been installed; need to check performance with Neon, possibly this weekend.
The absorber has been fitted and wrapped in MLI. Vacuum test offline, then remove endcaps and move on-line on Thursday and possibly start cool-down Friday.
Quench line is still in progress (not needed for Neon test) Walkabout with ISIS next Thu. but can't get final go-ahead until hardware complete.
Remote monitoring is in progress (see below).
Need to check the lead time for the Hydrogen gas delivery. (action)
AKurup: the remote monitoring NUC is in place and the network bridge has been configured. Need to set up the software IOC and then do GUI.
- Status of Support for Technical Endeavours And Knowledge (CWhyte)
No.
- AOB
- The VME IOCs are all reporting CPU temperatures of 126°C, which seems bogus; AKurup to contact DL about the readback. The system temperatures cover a range and have been up into the 40-50°C range, setting off warnings. Who is responsible for the crates and thus their temperature alarm thresholds? Daresbury. Ask e.g. Adrian, he'd know whom to pass it on to.
- CDB: Janusz has a fix that needs a restart of the CDB. AKurup will send an e-mail to the MICE list.
- DS controls (LHe level fluctuations): the DS SM got put into the manual state, which affects the scan rate of one of the PVs which then de-stabilised the control loop. Need to change the spreadsheet, other lessons to learn.
Updated by Nebrensky, Henry over 6 years ago · 10 revisions